A 67.3-kg man has a fever, and his average body temperature is 39.3oC. What mass of ice, initially at 0.0oC, should he consume in order to bring his average body temperature down to 36.3oC? Assume no heat loss to the surroundings, and neglect complexities introduced by metabolism

kg
Answer #1

cbody = specific heat of body = 3470

mbody = mass of body = 67.3 kg

L = latent heat of ice = 334 J /kg

cwater = specific heat of water = 4186

m = mass of ice

Heat gained by ice = heat lost by man

mL + m cwater (36.3 - 0 ) = mbody cbody (39.3 - 36.3)

m(334000) + m (4186) (36.7 - 0 ) = (67.3)(3470) (39.3 - 36.3)

m = 1.436 kg